The Ka of a monoprtic weak acid is 8.75 × 10^-3. What is the percent ionization of a 0.191 M solution of this acid?
Answer #1

HA dissociates as:

HA -----> H+ + A-

0.191 0 0

0.191-x x x

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((8.75*10^-3)*0.191) = 4.088*10^-2

since x is comparable c, our assumption is not correct

we need to solve this using Quadratic equation

Ka = x*x/(c-x)

8.75*10^-3 = x^2/(0.191-x)

1.671*10^-3 - 8.75*10^-3 *x = x^2

x^2 + 8.75*10^-3 *x-1.671*10^-3 = 0

This is quadratic equation (ax^2+bx+c=0)

a = 1

b = 8.75*10^-3

c = -1.671*10^-3

Roots can be found by

x = {-b + sqrt(b^2-4*a*c)}/2a

x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 6.762*10^-3

roots are :

x = 3.674*10^-2 and x = -4.549*10^-2

since x can't be negative, the possible value of x is

x = 3.674*10^-2

% dissociation = (x*100)/c

= 3.674*10^-2*100/0.191

= 19.2 %

Answer: 19.2 %